# Detecting Brittle Decisions for Free: Leveraging Margin Consistency in Deep Robust Classifiers

### v1

@@ Abstract
Despite extensive research on adversarial training strategies to improve robustness, the decisions of even the most robust deep learning models can still be quite sensitive to imperceptible perturbations, creating serious risks when deploying them for high-stakes real-world applications. While detecting such cases may be critical, evaluating a model’s vulnerability at a per-instance level using adversarial attacks is computationally too intensive and unsuitable for real-time deployment scenarios. The input space margin is the exact score to detect non-robust samples and is intractable for deep neural networks. This paper introduces the concept of margin consistency – a property that links the input space margins and the logit margins in robust models – for efficient detection of vulnerable samples. First, we establish that margin consistency is a necessary and sufficient condition to use a model’s logit margin as a score for identifying non-robust samples. Next, through comprehensive empirical analysis of various robustly trained models on CIFAR10 and CIFAR100 datasets, we show that they indicate strong margin consistency with a strong correlation between their input space margins and the logit margins. Then, we show that we can effectively use the logit margin to confidently detect brittle decisions with such models and accurately estimate robust accuracy on an arbitrarily large test set by estimating the input margins only on a small subset. Finally, we address cases where the model is not sufficiently margin-consistent by learning a pseudo-margin from the feature representation. Our findings highlight the potential of leveraging deep representations to efficiently assess adversarial vulnerability in deployment scenarios.
@@ 1 Introduction
Deep neural networks are known to be vulnerable to adversarial perturbations, visually insignificant changes in the input resulting in the so-called adversarial examples that alter the model’s prediction (Biggio et al., 2013; Goodfellow et al., 2015). They constitute actual threats in real-world scenarios (Evtimov et al., 2017; Gnanasambandam et al., 2021), jeopardizing their deployment in sensitive and safety-critical systems such as autonomous driving, aeronautics, and health care. Research in the field has been intense and produced various adversarial training strategies to defend against the vulnerability to adversarial perturbations with bounded $\ell_{p}$ norm (e.g., $p=2$, $p=\infty$) through augmentation, regularization, and detection (Xu et al., 2017; Madry et al., 2018; Zhang et al., 2019; Carmon et al., 2019; Wang et al., 2020; Wu et al., 2020; Rice et al., 2020), to cite a few. The empirical robustness (adversarial accuracy) of these adversarially trained models is still far behind their high performance in terms of accuracy. It is typically estimated by assessing the vulnerability of samples of a given test set using adversarial attacks (Carlini & Wagner, 2016; Madry et al., 2018) or an ensemble of attacks such as the standard AutoAttack (Croce & Hein, 2020b). The objective of that evaluation is to determine if, for a given normal sample, an adversarial instance exists within a given $\epsilon$-ball around it. Yet, this robustness evaluation over a specific test set gives a global property of the model but not a local property specific to a single instance (Seshia et al., 2018; Dreossi et al., 2019). Beyond that particular test set, obtaining this information for each new sample would typically involve rerunning adversarial attacks or performing a formal robustness verification, which in certain contexts may be computationally prohibitive in terms of resources and time. Indeed, in high-stakes deployment scenarios, knowing the vulnerability of single instances in real-time (i.e., their susceptibility to adversarial attacks) would be valuable, for example, to reduce risk, prioritize resources, or monitor operations. Current research lacks efficient and scalable ways to determine the vulnerability of a sample in a deployment context.
@@
The input space margin (i.e., the distance of the sample to the model’s decision boundary in the input space), or input margin in short, can be used as a score to determine whether the sample is non-robust and, as such, likely to be vulnerable to adversarial attacks. Computing the exact input margin is intractable for deep neural networks (Katz et al., 2017; Elsayed et al., 2018; Jordan & Dimakis, 2020). These input margins may not be meaningful for fragile models with zero adversarial accuracies as all samples are vulnerable (close to the decision boundary). However, for robustly trained models, where only certain instances are vulnerable, the input margin is very useful in identifying the critical samples. Previous research studies have explored input margins of deep neural networks during training, focusing on their temporal evolution (Mickisch et al., 2020; Xu et al., 2023), and their exploitation in improving adversarial robustness through instance-reweighting with approximations (Zhang et al., 2020; Liu et al., 2021) and margin maximization (Elsayed et al., 2018; Ding et al., 2020; Xu et al., 2023). However, to the best of our knowledge, no previous research studies the relationship between the input space margin and the logit margin of robustly trained deep classifiers in the context of vulnerability detection.
@@
Figure 1: Illustration of the input space margin, margin in the feature space and margin consistency. The margin-consistent model preserves the relative position of samples to the decision boundary in the input space to the feature space.
@@
  • We introduce the notion of margin consistency [^1] Code available at: https://github.com/ngnawejonas/margin-consistency, a property to characterize robust models that allow using their logit margin as a proxy estimation for the input space margin in the context of non-robust sample detection. We prove that margin consistency is a necessary and sufficient condition to reliably use the logit margin for detecting non-robust samples.
  • Through an extensive empirical investigation of pre-trained models on CIFAR10 and CIFAR100 with various adversarial training strategies, mainly taken from RobustBench (Croce et al., 2021), we provide evidence that almost all the investigated models display strong margin consistency, i.e., there is a strong correlation between the input margin and the logit margin.
  • We confirm experimentally that models with strong margin consistency perform well in detecting samples vulnerable to adversarial attacks based on their logit margin. In contrast, models with weaker margin consistency exhibit poorer performance. Leveraging margin consistency, we can also estimate the robust accuracy on an arbitrarily large test set by estimating the input margins only on a small subset.
  • For models where margin consistency does not hold, exhibiting a weak correlation between the input margin and the logit margin, we simulate margin consistency by learning to map the model’s feature representation to a pseudo-margin with a better correlation through a simple learning scheme.
(a) CIFAR10
(b) CIFAR100
Figure 2: Margin consistency of various models: there is a strong correlation between input space margin and logit margin for most $\ell_{\infty}$ robust models tested, the exceptions being DI0 and XU8 on CIFAR10. See Table 1 for the references on the models. The plots are given with standard error for the y-axis values in each interval.
@@ 2.1 Notation and Preliminaries
Notation We consider $f_{\theta}:\mathbb{R}^{n}\to\mathbb{R}^{C}$ a deep neural network classifier with weights $\mathbf{\theta}$ trained on a dataset of samples drawn iid from a distribution $\mathcal{D}$ on a product space $\mathcal{X}\times\mathcal{Y}$. Each sample $\mathbf{x}$ in the input space $\mathcal{X}\subset\mathbb{R}^{n}$ has a unique corresponding label $y\in\mathcal{Y}=\{1,2,\ldots,C\}$. The prediction of $\mathbf{x}$ is given by $\hat{y}(\mathbf{x})=\operatorname*{arg\,max}\nolimits_{j\in\mathcal{Y}}f_{\theta}^{j}(\mathbf{x})$, where $f_{\theta}^{j}(\mathbf{x})$ is the $j$-th component of $f_{\theta}(\mathbf{x})$. We consider that a deep neural classifier is composed of a feature extractor $h_{\psi}:\mathcal{X}\to\mathbb{R}^{m}$ and a linear head with $C$ linear classifiers $\{\mathbf{w_{j}},b_{j}\}$ such that $f_{\theta}^{j}(\mathbf{x})=\mathbf{w_{j}}^{\top}h_{\psi}(\mathbf{x})+b_{j}$. The predictive distribution $p_{\theta}(y|\mathbf{x})$ is obtained by taking the softmax of the output $f_{\theta}(\mathbf{x})$. A perturbed sample $\mathbf{x}^{\prime}$ can be obtained by adding a perturbation $\delta$ to $\mathbf{x}$ within an $\epsilon$-ball $B_{p}(\mathbf{x},\epsilon)$, an $\ell_{p}$-norm ball of radius $\epsilon>0$ centered at $\mathbf{x}$; $B_{p}(\mathbf{x},\epsilon):=\{\mathbf{x}^{\prime}:\|\mathbf{x}^{\prime}-\mathbf{x}\|_{p}=\|\delta\|_{p}<\epsilon\}$. The distance $\|\mathbf{x}^{\prime}-\mathbf{x}\|_{p}=\|\mathbf{\delta}\|_{p}$ represents the perturbation size defined as $\left(\sum\nolimits_{i=1}^{n}|\delta_{i}|^{p}\right)^{\frac{1}{p}}$. In this paper, we will focus on $\ell_{\infty}$ norm $\left(\|\mathbf{x}\|_{\infty}=\max\nolimits_{i=1,\ldots,n}|x_{i}|\right)$, which is the most commonly used norm in the literature.
@@
The logit margin is the difference between the two largest logits. For a sample $\mathbf{x}$ classified as $i=\hat{y}(\mathbf{x})=\operatorname*{arg\,max}\nolimits_{j\in\mathcal{Y}}f_{\theta}^{j}(\mathbf{x})$ the logit margin is defined as $\left(f_{\theta}^{i}(\mathbf{x})-\max\limits_{j,j\neq i}f_{\theta}^{j}(\mathbf{x})\right)>0$. It is an approximation of the distance to the decision boundary of $f_{\theta}$ in the feature space. The decision boundary in the feature space around $\mathbf{z}=h_{\psi}(\mathbf{x})$, the feature representation of $\mathbf{x}$, is composed of $(C-1)$ linear decision boundaries (hyperplanes) $\text{DB}_{ij}=\{\mathbf{z}^{\prime}\in\mathbb{R}^{m}:\mathbf{w}_{i}^{\top}\mathbf{z}^{\prime}+b_{i}=\mathbf{w}_{j}^{\top}\mathbf{z}^{\prime}+b_{j}\}$ ($j\neq i$). The margin in the feature space is, therefore, the distance to the closest hyperplane, i.e. $\min\limits_{j,j\neq i}d(\mathbf{z},\text{DB}_{ij})$, where the distance $d(\mathbf{z},\text{DB}_{ij})$ from $\mathbf{z}$ to a hyperplane $\text{DB}_{ij}$ has a closed-form expression:
@@
When the classifiers $\mathbf{w_{j}}$ are equidistant ($\|\mathbf{w}_{i}-\mathbf{w}_{j}\|_{q}=\omega>0,\,\forall i,j$), the margin becomes:
$\min\limits_{j,j\neq i}\frac{f_{\theta}^{i}(\mathbf{x})-f_{\theta}^{j}(\mathbf{x})}{\omega}=\frac{1}{\omega}\min\limits_{j,j\neq i}\bigg{(}f_{\theta}^{i}(\mathbf{x})-f_{\theta}^{j}(\mathbf{x})\bigg{)}=\frac{1}{\omega}\bigg{(}\underbrace{f_{\theta}^{i}(\mathbf{x})-\max\limits_{j,j\neq i}f_{\theta}^{j}(\mathbf{x})}_{\text{logit margin}}\bigg{)}.$ (3)
@@
$d_{out}(\mathbf{x})=f_{\theta}^{i}(\mathbf{x})-\max\limits_{j,j\neq i}f_{\theta}^{j}(\mathbf{x})$ (4)
@@ 2.2 Margin Consistency
We can evaluate margin consistency by computing the Kendall rank correlation ($\bm{\tau}\in[-1,1]$) between the output scores and the input margins over a test set. The Kendall rank correlation tests the existence and strength of a monotonic relationship between two variables. It makes no assumption on the distribution of the variables and is robust to outliers (Chattamvelli, 2024). Perfect margin consistency corresponds to an absolute value of $1$, and 00 means the absence of margin consistency.
@@ 2.3 Non-robust Samples Detection
$g(\mathbf{x})=\mathbbm{1}_{[d_{in}(\mathbf{x})\leq\epsilon]}(\mathbf{x})=\begin{cases}1&\text{ if }\,\,\mathbf{x}\,\,\text{ is non-robust}\\ 0&\text{ if }\,\,\mathbf{x}\,\,\text{ is robust}\\ \end{cases}.$
@@
Proof sketch. Fig. 3 presents intuition behind the proof of Theorem 1. For the first part of the theorem (see Fig. 3(a)), if there is a monotonic relationship between $d_{in}$ and $d_{out}$ (margin consistency), any point $\mathbf{x}$ with $d_{in}$ less than the threshold $\epsilon$ (non-robust) will also have $d_{out}$ less than $\lambda=d_{out}(\mathbf{x_{0}})$ (with $d_{in}(\mathbf{x_{0}})=\epsilon$). For the second part (see Fig. 3(b)), if there are two points $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ with non-concordant $d_{in}$ and $d_{out}$ (no margin consistency), then for a threshold $\epsilon_{0}$ between $d_{in}(x_{1})$ and $d_{out}(x_{2})$, they will both have different classes but no threshold of $d_{out}$ (horizontal line) can classify them both correctly. The complete proof of Theorem 1 is deferred to Appendix A.
@@
Common metrics for detection include (Hendrycks & Gimpel, 2017; Corbière et al., 2019; Zhu et al., 2023): the Area Under the Receiver Operating Curve (AUROC), which ensures the ability of a model to distinguish between the positive and negative classes across all possible thresholds; the Area Under the Precision-Recall Curve (AUPR), which evaluates the trade-off between precision and recall and is less sensitive to imbalance between positive and negative classes; and the False Positive Rate (FPR) at a 95% True Positive Rate (TPR) (FPR@95), that is crucial in systems where missing positive cases can have serious consequences, such as minimizing the number of vulnerable samples missed. The AUROC and AUPR of a perfect classifier is 1, while 0.5 for a random classifier.
2.4 Sample Efficient Robustness Evaluation
Margin consistency enables empirical robustness evaluation over an arbitrarily large test set by only estimating the input margins of a small subset of test samples. For a robustness evaluation at threshold $\epsilon$ (e.g., $\epsilon=8/255$ in $\ell_{\infty}$ norm on CIFAR10 and CIFAR100), we randomly sample a small subset of the large test set and determine the threshold $\lambda$ for the logit margin that corresponds to $\epsilon$. The threshold $\lambda$ is then used to detect vulnerable samples. With the true labels of these test sets, we can determine the proportion of correct non-vulnerable samples, which is the standard robust accuracy as described in Algorithm 1.
Algorithm 1 Sample Efficient Robustness Estimation
2:  Output: Robust Accuracy Estimation $\mathcal{A}_{r}$
3:  - Select uniformly at random a subset $X_{n}$ of $n$ samples from $X$.
A naive way to set the threshold $\lambda$ at line 6 of Algorithm 1 would be to set it to the detection threshold at $\alpha=95\%$ TPR or $\alpha=90\%$ TPR, but the logit margin threshold could vary from one model to ano; therefore a better way is to select it by tuning over values $\alpha>=0.80$ that gi ves the best approximation of the robust accuracy in terms of the absolute error on the small subset $X_{s}$. The same logic applies if we want to estimate the vulnerability of a large dataset without the labels.

### v3

@@ Abstract
Despite extensive research on adversarial training strategies to improve robustness, the decisions of even the most robust deep learning models can still be quite sensitive to imperceptible perturbations, creating serious risks when deploying them for high-stakes real-world applications. While detecting such cases may be critical, evaluating a model’s vulnerability at a per-instance level using adversarial attacks is computationally too intensive and unsuitable for real-time deployment scenarios. The input space margin is the exact score to detect non-robust samples and is intractable for deep neural networks. This paper introduces the concept of margin consistency – a property that links the input space margins and the logit margins in robust models – for efficient detection of vulnerable samples. First, we establish that margin consistency is a necessary and sufficient condition to use a model’s logit margin as a score for identifying non-robust samples. Next, through comprehensive empirical analysis of various robustly trained models on CIFAR10 and CIFAR100 datasets, we show that they indicate high margin consistency with a strong correlation between their input space margins and the logit margins. Then, we show that we can effectively and confidently use the logit margin to detect brittle decisions with such models. Finally, we address cases where the model is not sufficiently margin-consistent by learning a pseudo-margin from the feature representation. Our findings highlight the potential of leveraging deep representations to assess adversarial vulnerability in deployment scenarios efficiently.
@@ 1 Introduction
Deep neural networks are known to be vulnerable to adversarial perturbations, visually insignificant changes in the input resulting in the so-called adversarial examples that alter the model’s prediction (Biggio et al., 2013; Goodfellow et al., 2015). They constitute actual threats in real-world scenarios (Evtimov et al., 2017; Gnanasambandam et al., 2021), jeopardizing their deployment in sensitive and safety-critical systems such as autonomous driving, aeronautics, and health care. Research in the field has been intense and produced various adversarial training strategies to defend against the vulnerability to adversarial perturbations with bounded $\ell_{p}$ norm (e.g., $p=2$, $p=\infty$) through augmentation, regularization, and detection (Xu et al., 2017; Madry et al., 2018; Zhang et al., 2019; Carmon et al., 2019; Wang et al., 2020; Wu et al., 2020; Rice et al., 2020), to cite a few. The empirical robustness (adversarial accuracy) of these adversarially trained models is still far behind their high performance in terms of accuracy. It is typically estimated by assessing the vulnerability of samples of a given test set using adversarial attacks (Carlini & Wagner, 2016; Madry et al., 2018) or an ensemble of attacks such as the standard AutoAttack (Croce & Hein, 2020b). The objective of that evaluation is to determine if, for a given normal sample, an adversarial instance exists within a given $\epsilon$-ball around it. Yet, this robustness evaluation over a specific test set provides a global property of the model but not a local property specific to a single instance (Seshia et al., 2018; Dreossi et al., 2019). Beyond that specific test set, obtaining this information for each new sample would typically involve rerunning adversarial attacks or performing a formal robustness verification, which in certain contexts may be computationally prohibitive in terms of resources and time. Indeed, the computational cost makes it prohibitive to estimate robust accuracy at scale on large test sets and/or large models, for example, when using AutoAttack in standard mode. Moreover, in high-stakes deployment scenarios, knowing the vulnerability of single instances in real-time (i.e., their susceptibility to adversarial attacks) would be valuable, for example, to reduce risk, prioritize resources, or monitor operations. Therefore, there is a need for efficient and scalable ways to determine the vulnerability of a model’s decision on a given sample.
@@
The input space margin (i.e., the distance of the sample to the model’s decision boundary in the input space), or input margin in short, can be used as a score to determine whether the sample is non-robust and, as such, likely to be vulnerable to adversarial attacks. Computing the exact input margin is intractable for deep neural networks (Katz et al., 2017; Elsayed et al., 2018; Jordan & Dimakis, 2020). These input margins may not be meaningful for fragile models with zero adversarial accuracies as all samples are vulnerable (close to the decision boundary). However, for robustly trained models, where only certain instances are vulnerable, the input margin is very useful for identifying the critical samples. Previous research studies have explored input margins of deep neural networks during training, focusing on their temporal evolution (Mickisch et al., 2020; Xu et al., 2023), and their exploitation in improving adversarial robustness through instance-reweighting with approximations (Zhang et al., 2020; Liu et al., 2021) and margin maximization (Elsayed et al., 2018; Ding et al., 2020; Xu et al., 2023). However, to the best of our knowledge, no previous research studies the relationship between the input space margin and the logit margin of robustly trained deep classifiers in the context of vulnerability detection.
@@
Figure 1: Illustration of the input space margin, margin in the feature space and margin consistency. The model preserves the relative position of samples to the decision boundary in the input space to the feature space.
@@
  • We introduce the notion of margin consistency [^1] Code available at: https://github.com/ngnawejonas/margin-consistency, a property to characterize robust models that allow the use of their logit margin as a proxy estimation for the input space margin in the context of non-robust sample detection. We prove that margin consistency is a necessary and sufficient condition to reliably use the logit margin for detecting non-robust samples.
  • Through an extensive empirical investigation of pre-trained models on CIFAR10 and CIFAR100 with various adversarial training strategies, mainly taken from RobustBench (Croce et al., 2021), we provide evidence that almost all the investigated models display high margin consistency, i.e., there is a strong correlation between the input margin and the logit margin.
  • We confirm experimentally that models with high margin consistency perform well in detecting samples vulnerable to adversarial attacks based on their logit margin. In contrast, models with weaker margin consistency exhibit poorer performance.
  • For models where margin consistency does not hold, exhibiting a weak correlation between the input margins and the logit margins, we simulate margin consistency by learning to map the model’s feature representation to a pseudo-margin with a better correlation through a simple learning scheme.
@@ 2.1 Notation and Preliminaries
Notation We consider $f_{\theta}:\mathbb{R}^{n}\to\mathbb{R}^{K}$ a deep neural network classifier with weights $\mathbf{\theta}$ trained on a dataset of samples drawn iid from a distribution $\mathcal{D}$ on a product space $\mathcal{X}\times\mathcal{Y}$. Each sample $\mathbf{x}$ in the input space $\mathcal{X}\subset\mathbb{R}^{n}$ has a unique corresponding label $y\in\mathcal{Y}=\{1,2,\ldots,K\}$. The prediction of $\mathbf{x}$ is given by $\hat{y}(\mathbf{x})=\operatorname*{arg\,max}\nolimits_{k\in\mathcal{Y}}f_{\theta}^{k}(\mathbf{x})$, where $f_{\theta}^{k}(\mathbf{x})$ is the $k$-th component of $f_{\theta}(\mathbf{x})$. We consider that a deep neural network is composed of a feature extractor $h_{\psi}:\mathcal{X}\to\mathbb{R}^{m}$ and a linear head with $K$ linear classifiers $\{\mathbf{w_{k}},b_{k}\}$ such that $f_{\theta}^{k}(\mathbf{x})=\mathbf{w_{k}}^{\top}h_{\psi}(\mathbf{x})+b_{k}$. The predictive distribution $p_{\theta}(y|\mathbf{x})$ is obtained by taking the softmax of the output $f_{\theta}(\mathbf{x})$. A perturbed sample $\mathbf{x}^{\prime}$ can be obtained by adding a perturbation $\delta$ to $\mathbf{x}$ within an $\epsilon$-ball $B_{p}(\mathbf{x},\epsilon)$, an $\ell_{p}$-norm ball of radius $\epsilon>0$ centered at $\mathbf{x}$, $\{\mathbf{x}^{\prime}:\|\mathbf{x}^{\prime}-\mathbf{x}\|_{p}=\|\delta\|_{p}<\epsilon\}$. The distance $\|\mathbf{x}^{\prime}-\mathbf{x}\|_{p}=\|\mathbf{\delta}\|_{p}$ represents the perturbation size defined as $\left(\sum\nolimits_{i=1}^{n}|\delta_{i}|^{p}\right)^{\frac{1}{p}}$. In this paper, we will focus on $\ell_{\infty}$ norm $\left(\|\mathbf{x}\|_{\infty}=\max\nolimits_{i=1,\ldots,n}|x_{i}|\right)$, which is the most commonly used norm in the literature.
@@
The logit margin is the difference between the two largest logits. For a sample $\mathbf{x}$ classified as $i=\hat{y}(\mathbf{x})=\operatorname*{arg\,max}\nolimits_{j\in\mathcal{Y}}f_{\theta}^{j}(\mathbf{x})$ the logit margin is defined as $\left(f_{\theta}^{i}(\mathbf{x})-\max\limits_{j,j\neq i}f_{\theta}^{j}(\mathbf{x})\right)>0$. It is an approximation of the distance to the decision boundary of $f_{\theta}$ in the feature space. The decision boundary in the feature space around $\mathbf{z}=h_{\psi}(\mathbf{x})$, the feature representation of $\mathbf{x}$, is composed of $(K-1)$ linear decision boundaries (hyperplanes) $\text{DB}_{ij}=\{\mathbf{z}^{\prime}\in\mathbb{R}^{m}:\mathbf{w}_{i}^{\top}\mathbf{z}^{\prime}+b_{i}=\mathbf{w}_{j}^{\top}\mathbf{z}^{\prime}+b_{j}\}$ ($j\neq i$). The margin in the feature space is therefore the distance to the closest hyperplane $\min\limits_{j,j\neq i}d(\mathbf{z},\text{DB}_{ij})$, where the distance $d(\mathbf{z},\text{DB}_{ij})$ from $\mathbf{z}$ to a hyperplane $\text{DB}_{ij}$ has a closed-form expression:
@@
When the classifiers $\mathbf{w}_{j}$ are equidistant, i.e. $\|\mathbf{w}_{i}-\mathbf{w}_{j}\|_{q}=C$ whenever $i\neq j$, the margin becomes:
$\min\limits_{j,j\neq i}\frac{f_{\theta}^{i}(\mathbf{x})-f_{\theta}^{j}(\mathbf{x})}{C}=\frac{1}{C}\min\limits_{j,j\neq i}\bigg{(}f_{\theta}^{i}(\mathbf{x})-f_{\theta}^{j}(\mathbf{x})\bigg{)}=\frac{1}{C}\bigg{(}\underbrace{f_{\theta}^{i}(\mathbf{x})-\max\limits_{j,j\neq i}f_{\theta}^{j}(\mathbf{x})}_{\text{logitmargin}}\bigg{)}.$ (3)
@@
$d_{out}(\mathbf{x})=f_{\theta}^{i}(\mathbf{x})-\max\limits_{j,j\neq i}f_{\theta}^{j}(\mathbf{x}).$ (4)
@@ 2.2 Margin Consistency
We can evaluate margin consistency by computing the Kendall rank correlation ($\tau\in[-1,1]$) between the logit margins and the input margins over a test set. The Kendall rank correlation tests the existence and strength of a monotonic relationship between two variables. It makes no assumption on the distribution of the variables and is robust to outliers (Chattamvelli, 2024). While a positive value of $\tau$ indicates samples are ranked similarly (or identically for $\tau=1$) according to logit margins and input margins, a negative value of $\tau$ indicates that one margin’s ranking is roughly reversed. Perfect margin consistency corresponds to the situation $\tau=1$.
@@ 2.3 Non-robust Samples Detection
$g(\mathbf{x};f_{\theta})=\mathbbm{1}_{[d_{in}(\mathbf{x})\leq\epsilon]}=\begin{cases}1&\text{ if }\,\,\mathbf{x}\,\,\text{ is non-robust}\\ 0&\text{ if }\,\,\mathbf{x}\,\,\text{ is robust}\\ \end{cases}.$
@@
Proof sketch. Fig. 2 presents intuition behind the proof of Theorem 1. For the first part of the theorem (see Fig. 2(a)), if there is a monotonic relationship between $d_{in}$ and $d_{out}$ (margin consistency), any point $\mathbf{x}$ with $d_{in}$ less than the threshold $\epsilon$ (non-robust) will also have $d_{out}$ less than $\lambda_{0}=d_{out}(\mathbf{x_{0}})$ (with $d_{in}(\mathbf{x_{0}})=\epsilon$). For the second part (see Fig. 2(b)), if there are two points $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ with non-concordant $d_{in}$ and $d_{out}$ (no margin consistency), then for a threshold $\epsilon_{0}$ between $d_{in}(x_{1})$ and $d_{out}(x_{2})$, they will both have different classes but no threshold of $d_{out}$ can classify them both correctly. The complete proof of Theorem 1 is deferred to Appendix A. Common metrics for detection include (Hendrycks & Gimpel, 2017; Corbière et al., 2019; Zhu et al., 2023): the Area Under the Receiver Operating Curve (AUROC), which ensures the ability of a model to distinguish between the positive and negative classes across all possible thresholds; the Area Under the Precision-Recall Curve (AUPR), which evaluates the trade-off between precision and recall and is less sensitive to imbalance between positive and negative classes; and the False Positive Rate (FPR) at a 95% True Positive Rate (TPR) (FPR@95), that is crucial in systems where missing positive cases can have serious consequences, such as minimizing the number of vulnerable samples missed. The AUROC and AUPR of a perfect classifier is 1, while 0.5 for a random classifier.